# Through-Wall Imaging based on WiFi Channel State Information

### v1

@@ 1 Introduction
The capability of our method to facilitate visual monitoring in through-wall scenarios without conventional cameras is illustrated in Figure 1. In this example, the image in Figure 0(b) is reconstructed solely from the WiFi CSI amplitude spectrogram shown in Figure 0(c). Its close resemblance to the ground truth (Figure 0(a)) demonstrates the effectiveness of our approach and its potential for enhancing the interpretability of WiFi CSI by unlocking image-based downstream tasks. To stimulate further research in this emerging field, data, models, and code are made publicly available [^1] Supplementary Material, Zenodo (available from 27-10-2024).
@@ 3 Data
Our experimental setup, illustrated in Figure 2, is designed to jointly capture through-wall WiFi CSI and images of activities within a target room (3.8m×$\times$5.3m office). In this setup, a point-to-point WiFi transmitter-receiver arrangement is positioned outside the room, accompanied by an intra-room camera. The WiFi transmitter and receiver systems utilize the Espressif ESP32-S3-DevKitC-1U [^2] ESP32-S3-DevKitC-1U, https://docs.espressif.com/, (acc. 21-01-2024) development board, equipped with an ALFA Network APA-M25333APA-M25, https://alfa-network.eu/apa-m25, (acc. 21-01-2024) directional panel antenna featuring a ${}^{\circ}$ horizontal beam width and a gain of 8dBi @2.4GHz. The connection between transmitter and receiver is established using the Espressif ESP-NOW [^4] ESP-NOW, https://docs.espressif.com (acc. 21-01-2024) protocol, employing a fixed packet-sending rate of 100Hz. Images are captured at a resolution of 640$\times$480 pixels and a frame rate of 30Hz using an ESP32-S3-based camera board [^5] ESP32-S3-CAM, https://github.com/ (acc. 21-01-2024). To ensure temporal synchronization, both the WiFi receiver and camera are connected to a notebook that concurrently captures WiFi packets and images through serial over USB and WiFi connections, respectively.
@@ 4.3 Temporal Encoding
Model PSNRnormal-$\uparrow$ SSIMnormal-$\uparrow$ RMSEnormal-$\downarrow$ FIDnormal-$\downarrow$
UW
20.03$\pm$.05
0.734$\pm$.00
8.02$\pm$.04
142.95$\pm$2.4
GW
20.02$\pm$.07
0.734$\pm$.00
8.00$\pm$.05
141.00$\pm$5.0
C
20.39$\pm$.05
0.748$\pm$.00
7.83$\pm$.05
127.33$\pm$4.7
C+T 20.39$\pm$.04 0.749$\pm$.00 7.80$\pm$.02 125.62$\pm$3.2
@@ 5.1 Model Training
Utilizing the captured dataset, we train models based on the proposed MoPoE-VAE architecture for WiFi-CSI-based image synthesis, considering the following variants: uniform weighing (UW), Gaussian weighing (GW), and concatenation (C), as well as the combination of concatenation and temporal encoding (C+T). On top of this, a hyperparameter search is conducted, employing C+T on the validation subset, to determine suitable values for training hyperparameters. We evaluate over batch size $b\in$16, 32, 64, 128, 256, 512}, window size $L\in$51, 101, 151, 201, 251, 301}, and the $D_{KL}$ weighing parameter $\beta\in$1, 2, 4, 6}. This resulted in optimal values of $b=32$, $L=151$, and $\beta=1$, which we subsequently use for training all models. Concerning the recorded WiFi sequence, for each packet, we extract the 52 L-LTF subcarriers from the complex CSI matrix and precompute amplitudes which are used to sample $52\times L$ spectrograms for the CSI VAE to process. Furthermore, the raw images are rescaled to a resolution of 128$\times$128 pixels and normalized using the per-channel means and standard deviations of the dataset. For each configuration, utilizing the Adam optimizer ($lr=1e-3$), we conduct ten independent training runs spanning 50 epochs and select the model with the lowest validation loss for the final evaluation on the test subset.
Metrics. To assess the image reconstruction fidelity of our models, we employ metrics, well-established in compression/generative model research, such as the Peak Signal-to-Noise Ratio (PSNR), the Structural Similarity Index Measure (SSIM) [20], the Root Mean Squared Error (RMSE), and the Fréchet Inception Distance (FID) [21].
@@ 5.2 Quantitative Results
Table 1 summarizes the quantitative results of our ablation study, providing a comparison between the MoPoE-VAE aggregation options UW, GW, C, and C+T based on the PSNR, SSIM, RMSE, and FID metric, averaged over ten independent training runs. Among these models, UW and GW are the least effective, yielding comparatively low PSNR, SSIM, and high RMSE and FID scores. We improve upon this by introducing concatenation (C), which turns out to have a significant impact, evidenced by a higher PSNR and SSIM, along with lower RMSE and FID scores. This stark contrast in quantitative performance is also reflected visually, leading to notable enhancements in sharpness and perceptual fidelity. However, C+T (Figure 4(e)) outperforms all previous aggregation options, exhibiting slightly higher PSNR and SSIM, coupled with lower RMSE and FID scores, compared to C (Figure 4(d)). Additionally, we point out the observed improvement in FID scores for C+T, suggesting reduced perceptual distance and qualitative enhancements. Despite the only subtle impact of temporal encoding when combined with concatenation, reconstructed images are slightly sharper, and, more importantly, frame transitions are more coherent. In summary, while C+T demonstrates superior quantitative performance, its marginal differences highlight the importance of considering both quantitative metrics and visual assessments.
@@ 5.3 Qualitative Results
Our qualitative examination aligns with the trends observed in the quantitative analysis; Figure 5 illustrates the impact of different aggregation options on image reconstruction fidelity. In Figure 4(b), the uniform weighing (UW) model yields a blurry reconstruction, consistent with its lower PSNR and higher RMSE scores (see Table 1). We hypothesize that the lack of focus in the temporal domain results in diminished perceptual quality. Moving to Figure 4(c), the Gaussian weighting (GW) model showcases a slight improvement in quality despite the similarity in quantitative metrics. This improvement is attributed to the prioritization of central WiFi packets during weighing; that is, packets that are in closer temporal proximity to the ground truth image have a higher contribution, resulting in a clearer and more accurate reconstruction. Next, Figure 4(d) illustrates the concatenation (C) model: By learning the aggregation of WiFi features itself, not only further improves the sharpness of images, but even enhances the quality on a frame-to-frame level, eliminating spatiotemporal discontinuities, as illustrated in Figure 6. Quantitatively, C outperforms UW and GW, yet subtle artifacts and jittering remain visible in videos compared to C+T, possibly contributed by marginal differences in FID scores. Finally, the peak of our architectural improvements is reached when employing the C+T model, which is showcased in Figure 4(e), incorporating both concatenation and temporal encoding. Employing cyclic temporal encoding effectively mitigates the previously mentioned jittering artifact, resulting in smoother frame transitions and reduced FID scores. For a perspective on video quality, reconstructed videos complement these static images, showcasing for each aggregation option its ability to reconstruct images based on trough-wall WiFi CSI11footnotemark: 1.
@@
6 Future Work
While the MoPoE-VAE+C+T model achieves promising results on our dataset, effectively demonstrating the feasibility of image synthesis from through-wall WiFi CSI, we do not expect it to generalize to new people or environments without further adaptations. Generalization is, in general, an open problem in WiFi-based person-centric sensing, and achieving it across diverse scenarios in practice will require additional steps, posing a promising opportunity for future research [22].
In addition to improving generalization, another promising direction for future work involves expanding upon the modalities used in our approach. Employing the MoPoE-VAE framework, our model is inherently adaptable to multimodal data, leaving the flexibility for incorporating additional sensory data. For instance, combining WiFi CSI with radar, audio, or structural vibrations could provide complementary information and aid in scenarios where one modality alone may be insufficient for accurate reconstructions, e.g., in settings where the WiFi signal is disturbed or ambiguous due to environmental effects.
$i_{0}$
@@ 7 Conclusion
In this work, we have demonstrated the direct synthesis of person-centric images from WiFi CSI captured in a through-wall scenario. We collected a dataset of WiFi CSI and corresponding image time series, on the basis of which an ablation study was conducted to assess the fidelity of reconstructed images stemming from different architectural variations. From this evaluation, the model pairing the aggregation option concatenation with temporal encoding emerged as superior. Our results demonstrate the viability of the proposed approach, suggesting its potential for novel applications, such as through-wall visual monitoring without conventional cameras, and improving the interpretability of WiFi CSI by unlocking image-based downstream tasks, such as visual activity recognition.
@@ 8 Acknowledgments
This work is partially funded by the Vienna Business Agency (Blindsight-grant 4829418).

### v2

@@ 1 Introduction
The capability of our method to facilitate visual monitoring in through-wall scenarios without conventional cameras is illustrated in Figure 1. In this example, the image in Figure 1(b) is reconstructed solely from the WiFi CSI amplitude spectrogram shown in Figure 1(c). Its close resemblance to the ground truth (Figure 1(a)) demonstrates the effectiveness of our approach and its potential for enhancing the interpretability of WiFi CSI by unlocking image-based downstream tasks. To stimulate further research in this emerging field, data, models, and code are made publicly available [^1] Supplementary Material, https://github.com/StrohmayerJ/wificam.
@@ 3 Data
Our experimental setup, illustrated in Figure 2, is designed to jointly capture through-wall WiFi CSI and images of activities within a target room (3.8m×$\times$5.3m office). In this setup, a point-to-point WiFi transmitter-receiver arrangement is positioned outside the room, accompanied by an intra-room camera. The WiFi transmitter and receiver systems utilize the Espressif ESP32-S3-DevKitC-1U [^2] ESP32-S3-DevKitC-1U, https://docs.espressif.com/, (acc. 21-01-2024) development board, equipped with an ALFA Network APA-M25333APA-M25, https://alfa-network.eu/apa-m25, (acc. 21-01-2024) directional panel antenna featuring a 66 horizontal beam width and a gain of 8dBi @2.4GHz. The connection between transmitter and receiver is established using the Espressif ESP-NOW [^4] ESP-NOW, https://docs.espressif.com (acc. 21-01-2024) protocol, employing a fixed packet-sending rate of 100Hz. Images are captured at a resolution of 640$\times$480 pixels and a frame rate of 30Hz using an ESP32-S3-based camera board [^5] ESP32-S3-CAM, https://github.com/ (acc. 21-01-2024). To ensure temporal synchronization, both the WiFi receiver and camera are connected to a notebook that concurrently captures WiFi packets and images through serial over USB and WiFi connections, respectively.
@@ 5.1 Model Training
Utilizing the captured dataset, we train models based on the proposed MoPoE-VAE architecture for WiFi-CSI-based image synthesis, considering the following variants: uniform weighing (UW), Gaussian weighing (GW), and concatenation (C), as well as the combination of concatenation and temporal encoding (C+T).
On top of this, a hyperparameter search is conducted, employing C+T on the validation subset, to determine suitable values for training hyperparameters. We evaluate over batch size $b\in\{16,32,64,128,256,512\}$, window size $L\in\{51,101,151,201,251,301\}$, and the $D_{KL}$ weighing parameter $\beta\in\{1,2,4,6\}$. This resulted in optimal values of $b=32$, $L=151$, and $\beta=1$, which we subsequently use for training all models.
Concerning the recorded WiFi sequence, for each packet, we extract the 52 L-LTF subcarriers from the complex CSI matrix and precompute amplitudes used to sample $52\times L$ spectrograms for the CSI VAE to process. Furthermore, the raw images are rescaled to a resolution of 128$\times$128 pixels and normalized using the per-channel means and standard deviations of the dataset. For each configuration, utilizing the Adam optimizer ($lr=1e-3$), we conduct ten independent training runs spanning 50 epochs and select the model with the lowest validation loss for the final evaluation on the test subset.
Metrics. To assess the image reconstruction fidelity of our models, we employ metrics well-established in compression/generative model research, such as the Peak Signal-to-Noise Ratio (PSNR), the Structural Similarity Index Measure (SSIM) [20], the Root Mean Squared Error (RMSE), and the Fréchet Inception Distance (FID) [21].
[…]
UW 20.03$\pm$.05 0.734$\pm$.00 8.02$\pm$.04 142.95$\pm$2.4
GW 20.02$\pm$.07 0.734$\pm$.00 8.00$\pm$.05 141.00$\pm$5.0
C 20.39$\pm$.05 0.748$\pm$.00 7.83$\pm$.05 127.33$\pm$4.7
C+T 20.39$\pm$.04 0.749$\pm$.00 7.80$\pm$.02 125.62$\pm$3.2
@@ 5.2 Quantitative Results
Table 1 summarizes the quantitative results of our ablation study, providing a comparison between the MoPoE-VAE aggregation options UW, GW, C, and C+T based on the PSNR, SSIM, RMSE, and FID metric, averaged over ten independent training runs. Among these models, UW and GW are the least effective, yielding comparatively low PSNR, SSIM, and high RMSE and FID scores.
We improve upon this by introducing concatenation (C), which turns out to have a significant impact, evidenced by a higher PSNR and SSIM, along with lower RMSE and FID scores. This stark contrast in quantitative performance is also reflected visually, leading to notable enhancements in sharpness and perceptual fidelity. However, C+T (Figure 5(e)) outperforms all previous aggregation options, exhibiting slightly higher PSNR and SSIM, coupled with lower RMSE and FID scores, compared to C (Figure 5(d)). Additionally, we point out the observed improvement in FID scores for C+T, suggesting reduced perceptual distance and qualitative enhancements.
Despite the only subtle impact of temporal encoding when combined with concatenation, reconstructed images are slightly sharper, and, more importantly, frame transitions are more coherent. In summary, while C+T demonstrates superior quantitative performance, its marginal differences highlight the importance of considering both quantitative metrics and visual assessments.
@@ 5.3 Qualitative Results
Our qualitative examination aligns with the trends observed in the quantitative analysis; Figure 5 illustrates the impact of different aggregation options on image reconstruction fidelity. In Figure 5(b), the uniform weighing (UW) model yields a blurry reconstruction, consistent with its lower PSNR and higher RMSE scores (see Table 1). We hypothesize that the lack of focus in the temporal domain results in diminished perceptual quality. Moving to Figure 5(c), the Gaussian weighting (GW) model showcases a slight improvement in quality despite the similarity in quantitative metrics. This improvement is attributed to the prioritization of central WiFi packets during weighing; that is, packets that are in closer temporal proximity to the ground truth image have a higher contribution, resulting in a clearer and more accurate reconstruction. Next, Figure 5(d) illustrates the concatenation (C) model: By learning the aggregation of WiFi features itself, not only further improves the sharpness of images but even enhances the quality on a frame-to-frame level, eliminating spatiotemporal discontinuities, as illustrated in Figure 6. Quantitatively, C outperforms UW and GW, yet subtle artifacts and jittering remain visible in videos compared to C+T, possibly contributed by marginal differences in FID scores. Finally, the peak of our architectural improvements is reached when employing the C+T model, which is showcased in Figure 5(e), incorporating both concatenation and temporal encoding. Employing cyclic temporal encoding effectively mitigates the previously mentioned jittering artifact, resulting in smoother frame transitions and reduced FID scores. For a perspective on video quality, reconstructed videos complement these static images, showcasing for each aggregation option its ability to reconstruct images based on through-wall WiFi CSI11footnotemark: 1.
@@
6 Limitations
While the MoPoE-VAE+C+T model shows promising results on our dataset, demonstrating the feasibility of image synthesis from through-wall WiFi CSI, it may not generalize to new people or environments without further adaptations. Generalization is an open problem in WiFi-based person-centric sensing, requiring additional steps across diverse scenarios [22]. Enhancing cross-domain generalization could involve domain-invariant feature extraction techniques or training data from diverse domains without needing architectural changes. Additionally, capturing samples in the target domain for fine-tuning or test-time training could improve generalization in practical applications. Another promising direction for future work is expanding the modalities used in our approach. The MoPoE-VAE framework is adaptable to multimodal data, allowing for additional sensory inputs. Combining WiFi CSI with radar, audio, or structural vibrations could provide complementary information and help in scenarios where one modality alone is insufficient for accurate reconstructions, such as when the WiFi signal is disturbed or ambiguous due to environmental effects.
$i_{1}$
@@ 7 Conclusion
In this work, we have demonstrated the direct synthesis of person-centric images from WiFi CSI captured in a through-wall scenario. We collected a dataset of WiFi CSI and corresponding image time series, based on which an ablation study was conducted to assess the fidelity of reconstructed images stemming from different architectural variations. From this evaluation, a model pairing with the aggregation option concatenation and temporal encoding emerged as superior. Our results demonstrate the viability of the proposed approach, suggesting its potential for novel applications, such as through-wall visual monitoring without conventional cameras, and improving the interpretability of WiFi CSI by unlocking image-based downstream tasks, such as visual activity recognition.
@@ 8 Acknowledgments
This work is partly funded by the Vienna Business Agency (grant 4829418) and the Austrian security research program KIRAS of the Austrian Research Promotion Agency FFG (grant 49450173).